### v5

@@ Abstract
We tackle the hard problem of consciousness taking the naturally selected, embodied organism as our starting point. We provide a formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence. Such interpretations imply behavioural policies which are differentiated from each other only by the qualitative aspect of information processing. Natural selection favours systems that intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate interventions. This produces interoceptive and exteroceptive classifiers and determines priorities. In formalising the seminal distinction between access and phenomenal consciousness, we claim that access consciousness at the human level requires the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others, and that this requires phenomenal consciousness. Phenomenal without access consciousness is likely common, but the reverse is implausible. To put it provocatively: death grounds meaning, and Nature does not like zombies. We then describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies. Our proposal lays the foundation of a formal science of consciousness, closer to human fact than zombie fiction.
keywords:
hard problem of consciousness
@@ 1 Introduction
Why is anything conscious? Both biological and other physical [^1] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet humans consciously experience as well as process information. Why? Living organisms are constantly processing self and world related information to survive in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space (e.g. occupy space). Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [1, 2].
Initially formally introduced in the field of cybernetics [3, 4] the notion of self-organisation has been applied to various disciplines including physics [5], biology [6, 7] and neuroscience [8, 9, 10]. Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [11, 12, 13].
@@
This question is the subject of long-standing debate [14]. One highly influential view is that consciousness has two aspects [14]. The first is functional, by which we mean the ability to access and communicate information [15]. How function relates to consciousness is considered the “easy problem” of consciousness [2]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [2, 15, 1, 16, 17]. This doesn’t mean just global states like being awake, but more local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [14]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [15]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
The hard problem has sparked a substantial body of work, detailed discussion of which [14, 18] lies beyond the scope of our paper.
In what follows we start with the observation that the mental/physical distinction that fuels the classical debates, ranging from panpsychism [19] to physicalism à la Dennett [20] may not be the only way to tackling the hard problem of consciousness. We remain agnostic to the question whether conceivability arguments such as the zombie thought experiment [21] or the knowledge argument [22, 23] are the best way to approach the problem of phenomenal consciousness. For example, one may argue that artificial systems such as Large Language Models (LLMs) conceive entities that humans never conceived or never will. It remains an open question on whether these entities exist just on the basis that a system can conceive them abstractedly or not. Our aim is more modest and starts with low level biological facts rather than high level philosophical abstractions to reformulate the problem of consciousness in a way that does not take for granted the mental/physical division in the first place. In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
@@
Take for example the higher order thought theory (HOT) [24, 25] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
HOTs are a good starting point to understand why most biological information processing goes on “in the dark”, but why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? Is there something it is like to “be in the dark” processing information at the lower bodily levels?
@@
Longtime considered a fringe approach, the embodied cognition paradigm [26] has recently gained substantial influence in cognitive science and philosophy [27, 28, 29]. The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
If this is so then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order levels of information processing only. In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
@@
Our approach is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [30] based on Global Workspace Theory [31], and constructivist approaches to artificial intelligence [32] take similar position. It is our explanation of how and why that is novel.. We establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark”.
@@
Enactivism is roughly the view that information processing arises through interaction between an organism and its environment [33], and it is considered by some to be incompatible with narrow definitions of computation. The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [34, 35]. Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another, so it is compatible with enactivism. Pancomputationalism is the idea that all systems are computational. We build upon a formalism called pancomputational enactivism [36].
@@
To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [37] - we are applying Sutton’s “bitter lesson” to biology.. First, in Section 2 we integrate and extend separately published works of narrower scope [36, 38, 39] to justify our model of self-organising systems. In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories. Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence. This takes us from attraction and repulsion to physical states, through lower order thoughts to higher order meta representations. We call this the psychophysical principle of causality. In Section 5 we extend previous work [39] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures. In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self. We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible. In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures, with examples from nematodes to humans.
@@ 2 Back to Foundations
To meet length restrictions, the full mathematical definitions have been placed in the appendices [40]. The core arguments should be understandable without the math, but for convenience here is a quick reference guide:
@@
    A set $\Phi$ is assumed, whose elements are called states. A declarative program is set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$. $P$ is the set of all programs. Embodiment is formalised as an abstraction layer, for which we single out finite set $\mathfrak{v}\in 2^{\Phi}$ called a vocabulary. A choice of vocabulary $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$.
@@
    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcup l$. A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t. $l_{1}\subseteq l_{2}$. The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t. $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. The extension of a set $X\subseteq{L_{\mathfrak{v}}}$ of statements is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
@@
    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the outputs available given ${I_{\alpha}}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs. All the tasks an abstraction layer tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$. A parent is higher level than its children.
@@
    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$. Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}$. A task is complete if $o\in O_{\alpha}$. Given a task definition $\alpha$ s.t. $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
@@
    A causal identity for $\mathfrak{b}$ is a policy classifying the causal interventions undertaken by $\mathfrak{b}$. Subscript denotes the system that learned the causal identity, while superscript denotes the object it classifies. For example, $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity learned by $\mathfrak{a}$ to classify $\mathfrak{b}$’s interventions, and $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
We begin by formalising all conceivable environments [36, 41] from first principles:
@@
Hence, the environment is a set $\Phi$ of mutually exclusive states. We call a set of states a declarative “program” [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [42].. A program is “true” relative to the states it contains. For example, if time is one dimension along which things change, then there is one state at a time. A “fact” is a program which contains the state at the present time. It follows that every aspect of the environment is a set of programs. An aspect of the environment is realised by (exists in) a state if the programs it contains are facts. Previous work argues this formalises every conceivable environment [41].
To give some intuition as to why, note that we assume nothing of what states might be or contain. We assume no differences within a state, only between states, and those differences are the programs of which aspects are formed. After all, one can only interact with aspects of one’s environment. If one were to try and pinpoint what an aspect was made of, the answer would be deferred to other aspects. Any description that sought to fully encapsulate the semantics, truth or unmediated pure experience of a state would always be delayed or incomplete. This does not render such attempts vacuous, but speaks to their inherent contingency and conditionality, which is what this formalism attempts to capture.

### v6

@@ Abstract
We tackle the problem of consciousness by taking the naturally selected, embodied organism as our starting point. We provide a formalism describing how biological systems such as human bodies self-organize to hierarchically interpret unlabelled sensory information according to valence. The system is attracted and repelled at different spatial and temporal scales. This is a qualitative interpretation of an unlabelled physical state. We show how such interpretations imply behavioural policies which are differentiated from each other only by this qualitative aspect of information processing. Natural selection favours systems that actively intervene in the world to achieve homeostatic and reproductive goals. Put provocatively, death grounds meaning. This means that in living systems information processing is necessarily subjective, that is, it has quality embedded into its very core. Qualitative information processing involves interoceptive and exteroceptive classifiers, and determines priorities for self-survival. We formulate The Psychophysical Principle of Causality as a theorem, and prove generalisation optimal learning forces this valence first ontology. Qualitative good or bad processing necessarily comes before quality neutral representations of properties (i.e. “red” is constructed from valence). Under selection pressures like sophisticated predation this produces a hierarchy of selves, of which reafference and reflective self awareness are a consequence. We discuss this in light of the seminal distinction between phenomenal and access consciousness. We claim that phenomenal consciousness without access is likely common, but the reverse is implausible. Our proposal lays the foundation of a formal science of consciousness, closer to human fact than zombie fiction.
@@ 1 Introduction
Why is anything conscious? Both biological and other physical systems process information, yet humans consciously experience as well as process the information. Why? Living organisms are constantly processing self and world related information to survive in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space (e.g. occupy space). Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation Nagel (1974); Chalmers (1995).
Initially formally introduced in the field of cybernetics Ashby (1947); von Foerster (1960) the notion of self-organisation has been applied to various disciplines including physics Haken (1983), biology Camazine (2003); Bell and Deater-Deckard (2007) and neuroscience Kelso (1995); Friston (2010); Tognoli and Kelso (2014). Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment Camazine et al. (2001); Seeley (2002); Rosas et al. (2018).
@@
This question is the subject of long-standing debate Seth and Bayne (2022). One highly influential view is that consciousness has two aspects Seth and Bayne (2022). The first is functional, by which we mean the ability to access and communicate information Block (1995). How function relates to consciousness is considered the “easy problem” of consciousness Chalmers (1995). The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness Chalmers (1995); Block (1995); Nagel (1974); Gallagher and Zahavi (2021); Fuchs (2017). This doesn’t mean just global states like being awake, but more local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them Seth and Bayne (2022). It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia Block (1995). For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is somewhat related to the “hard problem” of consciousness. However the hard problem has sparked a substantial body of work, detailed discussion of which Seth and Bayne (2022); Northoff (2014) lies beyond the scope of our paper. In what follows we largely avoid the mental/physical distinction that fuels the classical debates, ranging from panpsychism Goff (2019) to physicalism à la Dennett Dennett (1991).
Likewise we remain agnostic to the question whether conceivability arguments such as the zombie thought experiment Chalmers (1996) or the knowledge argument Jackson (1982, 1986) are the best way to approach the problem of phenomenal consciousness.
How we do so can be understood in terms of the hardware and software metaphor. The distinction between mental and physical that fuels classical debates can be understood in computational terms. In this view mental activity is like software, which interacts with the physical world through hardware Hutter (2010); Bennett (2025c); Orseau and Ring (2012). Hardware acts as an interpreter between software and the physical environment. This is somewhat like the Pineal gland in Cartesian dualism, and so this view is called “computational dualism” Bennett (2024b). While software is a useful frame for building standardised products, formalisms of intelligence based on computational dualism are inherently flawed Bennett (2024b); Leike and Hutter (2015). The problem is that software is not in fact a separate substance. In actuality, software is a state of the hardware on which it runs. To explain, computers can be understood as “stacks” of abstraction layers, where each layer is simplified subsets of behaviour of the layer below. For example Python code relies on interpretation by a program written in C, which is built on Assembly, which in turn relies on machine code and thus hardware. In this sense, software is an “abstraction layer” over the hardware on which it runs. Nature privileges adaptation, and working only at high levels of abstraction can limit adaptability, decreasing efficiency Bennett (2025a). Compare the vast quantity of data and energy required to train a large language model to solve a problem, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they delegate adaptation to lower levels of abstraction Bennett (2024a). For example, an organism can be understood as an abstraction layer over organs, which are a layer over cells. Adaptation takes place at every level of this biological “stack”, whereas a software “intelligence” in a computer is adaptive only at a a high level of abstraction, in software. Our ability to model and understand problems is limited by the level of abstraction at which we view them Bennett (2025a). Efficiency problems in Artificial Intelligence (AI) can be addressed by accounting for abstraction layers Bennett (2024b). Mental activity can be framed as a layer over the body, so rather than trying to explain the mind in the abstract like software, we start at the level of the embodied organism.
In short, our explanation of consciousness asserts that world has valence. Physical states are attractive or repulsive to a self-organising adaptive system. Organisms are such self-organising systems which learn to classify causes of valence that affect their decision procedures, to have a better chance of surviving. Qualia and affects, the atomic building blocks of phenomenal consciousness, are qualitative interpretations rather than abstract logical or mathematical representations. Qualitative processing conveys an advantage in evolutionary terms. At high levels of abstraction these appear to be irreducible qualities to which valence seems attached. However at lower levels of abstraction these qualities are the various states of attraction and repulsion happening in a body at a given point in time. Put another way, sensory stimuli are attractive or repulsive at multiple spatial and temporal scales within a self-organising system Bennett (2025a). Qualia and affect are easily and obviously linked to affordances in the world by organisms. Phenomenology has a ‘logic’ – not the symbols and rules of the supersets of first order logic that make up programming languages but a presymbolic ‘logic’ of cause and effect. These are modelled in the decision procedures of prelinguistic organisms as cause and affect. At a high level, organic decision procedures boil down to affect-infused qualia, the key components of phenomenology. Natural selection favours organisms with phenomenology. Not always but often. Our position is that phenomenology is functional and decisive, not epiphenomenal or illusory, but this is not to deny that conscious decisions have subconscious components. We see consciousness as the “executive” top level of a “stack” of biological components.
We bypass the “computational dualism” of hardware and software Bennett (2024b) by starting with low level biological facts rather than high level philosophical abstractions to reformulate the problem of consciousness in a way that does not take for granted the mental/physical division in the first place.
@@
Take for example the higher order thought theory (HOT) Rosenthal (2005); Brown et al. (2019) holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness. HOTs might be a good starting point to understand why most biological information processing goes on “in the dark”, but why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? Is there something it is like to “be in the dark” processing information at the lower bodily levels?
@@
Long considered a fringe approach, the embodied cognition paradigm Varela et al. (2016) has recently gained substantial influence in cognitive science and philosophy Ciaunica and others (2023); McMillen and Levin (2024); Seth and Tsakiris (2018). The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
If this is so then unravelling consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order, cognitive levels of information processing. In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system. Rather they fundamentally constitute what a conscious experience is. How can we prove that?
@@
Our aim here is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the problem by showing the phenomenal to be intrinsically functional [^1] Note that we aren’t the first to claim that the phenomenal is functional. For example, the Conscious Turing Machine Blum and Blum (2020) based on Global Workspace Theory Baars (1997), and constructivist approaches to AI Wang (2020) take similar positions. It is our explanation of how and why that is novel.. We establish axioms that hold in every conceivable environment. We then argue that if those axioms hold there is no conceivable environment where the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark” [^2] That is, without explicit conscious awareness of it..
Box 1: Supporting formal results and provenance This paper is primarily philosophical. However, the key claims are also supported by formal theorems, proofs and experiments provided in the Supplementary Information (ESM). Many of these experiments and proofs were peer reviewed and published in the PhD thesis “How To Build Conscious Machines” Bennett (2025a), and in the proceedings of the Spring 2026 AAAI Workshop on Machine Consciousness Bennett (2026b). The main manuscript also introduces new formal results. These include the proof of The Psychophysical Principle of Causality and substantially improved propositions showing how the 1ST, 2ND and 3RD order selves follow from generalisation optimal learning under the scale and incentive preconditions. Full proofs are included in the ESM. The reformulation of informal The Psychophysical Principle of Causality as a formal theorem, and the resulting proof, are dependent upon the already published theorem The Law of the Stack Bennett (2024a), so The Law of the Stack is also detailed in the ESM. Boxes 2 to 4 summarise the new theorem, the other proof results, and the experiments.
@@
Enactivism is roughly the view that information processing arises through interaction between an organism and its environment Thompson (2007), and it is considered by some to be incompatible with narrow definitions of computation. The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression Piccinini (2015); Piccinini and Maley (2021). Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another. To formalise enactivism [^3] By formalising something akin to “intelligence without representation” as described by Brooks Brooks (1991), we end up with physical interpretations rather than representations., we build upon Pancomputational Enactivism Bennett (2024b, a) and Stack Theory Bennett (2025a), which frame intelligence in terms of embodied abstraction layers.
@@
To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^4] This is loosely inspired by a scale-based framing of machine learning Sutton (2019) - we are applying Sutton’s “bitter lesson” to biology.. First, in Section 2 we integrate and extend separately published works of narrower scope Bennett (2024b, 2023c, 2023a) to justify our model of self-organising systems. In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories. Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence. This takes us from attraction and repulsion to physical states, through lower order thoughts to higher order meta representations. In the associated thesis Bennett (2025a) we call this the psychophysical principle of causality. Here we formalise this as a theorem of the same name (theorem 3 in the ESM). In Section 5 we extend previous work Bennett (2023a) on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures. In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self. We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible. In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures, with examples from nematodes to humans.
@@ 2 Back to Foundations
This paper can be read in purely narrative form, meaning the core argument should be understandable without the math. However for readers who do wish to delve into the mathematical details, a short summary of definitions is given below. The full set of definitions used in this paper, along with full proofs and the full experimental protocols and plots, is in the Electronic Supplementary Material (ESM). The longer Stack Theory appendix Bennett (2025a) is the master reference.
@@
    An environment is a nonempty set $\Phi$ of mutually exclusive states. A (declarative) program is any set of states $p\subseteq\Phi$. Let $\mathcal{P}:=2^{\Phi}$ be the set of all programs. An abstraction layer is a finite vocabulary $\mathfrak{v}\subseteq\mathcal{P}$, meaning the finite set of programs a body can stably implement. A vocabulary induces an embodied formal language
    $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap_{p\in l}p\neq\emptyset\}.$
@@
    Members of $L_{\mathfrak{v}}$ are called statements. The truth set of a statement $l$ is
    $T(l):=\bigcap_{p\in l}p.$
    We take $T(\emptyset)=\Phi$ by convention. The statement $l$ is true relative to a state $\phi\in\Phi$ iff $\phi\in T(l)$. A completion of $x$ is any statement $y$ such that $x\subseteq y$. The extension of a statement $x$ is the set of all its completions
    $E_{x}:=\{y\in L_{\mathfrak{v}}:x\subseteq y\}.$
    For a set $X\subseteq L_{\mathfrak{v}}$ we write $E_{X}:=\bigcup_{x\in X}E_{x}$. Two statements are equivalent when they have the same extension.
@@
    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subseteq L_{\mathfrak{v}}$ are the inputs and $O_{\alpha}\subseteq E_{I_{\alpha}}$ are the correct outputs. Tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
@@
    A policy is a statement $\pi\in L_{\mathfrak{v}}$. A policy is correct for $\alpha$ iff
    $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}.$
    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose an output $o\in E_{i}\cap E_{\pi}$. A task is complete if $o\in O_{\alpha}$.
@@
    Learning must choose among many correct policies. A standard proxy is weakness $w(\pi):=|E_{\pi}|$, which counts how many further commitments remain compatible with $\pi$ while staying correct. Weakness maximisation (w-maxing) chooses a correct policy that maximises $w(\pi)$. A causal identity is a policy $c^{\mathfrak{b}}_{\mathfrak{a}}\in L_{\mathfrak{v}}$ that classifies interventions by some source $\mathfrak{b}$ as distinct from matched observations, relative to a learner $\mathfrak{a}$. Nested causal identities represent higher orders of modelling. For example $c^{\mathfrak{ba}}_{\mathfrak{a}}\subseteq c^{\mathfrak{b}}_{\mathfrak{a}}$ is a refinement of $\mathfrak{a}$’s model of $\mathfrak{b}$ that depends on $\mathfrak{a}$’s prediction of how $\mathfrak{b}$ models $\mathfrak{a}$.
We begin by formalising all conceivable environments Bennett (2024b, c) from first principles:
@@
Hence, the environment is a set $\Phi$ of mutually exclusive states. We call a set of states a declarative “program” [^5] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative Howard (1980).. A program is “true” relative to the states it contains. For example, if time is one dimension along which things change, then there is one state at a time. A “fact” is a program which contains the state at the present time. It follows that every aspect of the environment is a set of programs. An aspect of the environment is realised by (exists in) a state if the programs it contains are facts. Given each state is a difference, it follows that states are mutually exclusive, which elsewhere has been used to portray time as difference and each state as a point in time [^6] Each possible sequence of states might be thought of as a timeline, and the set of all timelines can be thought of as the set of all conceivable universes (infinite if the set of states is infinite).. Such foundational arguments are largely beyond the scope of this paper and are discussed at length elsewhere Bennett (2025a). However, this frame allows us to talk about “computation” in a much more general sense than Turing computation. By computation, we mean physical cause and effect, dictated by the state transitions of whatever environment we might conceive of. In the frame of Stack Theory Bennett (2025a, 2024a), this is like saying we don’t know where the base abstraction layer is, whether there even is such a layer, or what it might entail. We are merely saying that no matter what abstraction layer is considered, there must be change or difference for there to be anything at all. That entails there must be states, which entails there are ontological “facts” or “true programs” we can formalise as sets of states, every aspect of the environment that exists is a set of such facts, and finally this lets us talk about computation in a much more general sense than Turing computation [^7] Related work holds each timeline is a “cosmic ought” from which goals and value judgements follow Bennett (2025a). From a subjective perspective of a body (an aspect of the environment), “truth” is unknown (there can be many possible timelines in which a given body exists, seeming non-deterministic), but from the objective perspective of a given timeline truth is deterministic. Deterministic and non-deterministic views can be reconciled, just as the set of all languages recognisable as DFAs (deterministic finite automata) is the same as that recognisable as NFAs (non-deterministic) Rabin and Scott (1959). In this view, the universe has a sort of “opinion” in the sense that a given timeline preserves some aspects of the environment and destroys others as it progresses. The set of all possible timelines would be the set of all such “opinions”.. We can talk about any given environment, or any given physical body within that environment, as an abstraction layer that performs a unique set of computations [^8] Put another way, no two bodies perform exactly the same set of computations.. To give some more intuition as to why we can talk about any body, or any conceivable environment in this frame, note that we assume nothing of what states might ontologically be or epistemologically contain. We assume no differences within a state, only between states, and those differences are the programs of which aspects are formed. After all, one can only interact with aspects of one’s environment. If one were to try and pinpoint what an aspect was made of, the answer would be deferred to other aspects. Any description that sought to fully encapsulate the semantics, truth or unmediated pure experience of a state would always be delayed or incomplete. This does not render such attempts vacuous, but speaks to their inherent contingency and conditionality, which is what this formalism attempts to capture.